how much money has to be invested at 2.9% interest compounded continuously to have $34000 after 18 years?

Answer by lwsshak3(11628)

how much money has to be invested at 2.9% interest compounded continuously to have $34000 after 18 years?
***
formula for continuous compounding: A=Pe^rt, P=initial investment, r=interest rate, t=years, A=amt after t-years.
For given problem:
P=?
r=.029
t=18
A=$34000
..
A=Pe^rt
P=A/e^rt=34000/e^(.029*18)=34000/e^(0.522)=34000/1.6854=20173
Amount of money has to be invested=$20,173